# Web-CogReasoner: Towards Knowledge-Induced Cognitive Reasoning for Web Agents

### v1

@@ 4.2 Web-CogReasoner Framework
      Review the updated list of Apple Stores;
    3. 3.
      (Fallback) Click the “Complete store list” if results are missing.
    This layer translates semantic understanding into executable strategies—effectively answering “How should the task be accomplished?”
By modularizing the CoT along these knowledge axes, Web-CogReasoner achieves structured planning grounded in a cognitively coherent framework. Unlike conventional end-to-end models that produce opaque or hallucinated justifications, each reasoning step here is explicitly traceable to factual evidence, semantic inference, or procedural logic. This alignment reduces ungrounded output and enhances interpretability.
As shown in Figure 4, the color-coded CoT segments (yellow, purple, blue) illustrate how the agent transitions from perception to comprehension to action. This workflow encapsulates the chain:
[…]
Through this process, the Web-CogReasoner becomes a practical embodiment of Web-CogKnowledge, leveraging its hierarchical structure to support robust, transparent, and goal-directed web reasoning.
4.3 Training
We employ a multi-stage Imitation Learning strategy to train our model on the Web-CogDataset, utilizing Qwen2.5-VL-7B bai2025qwen2 as the base model. Each stage is aligned with a distinct layer of the Web-CogKnowledge Framework: (1) the first knowledge content learning focuses on acquiring Factual Knowledge and Conceptual Knowledge, enabling the model to interpret web content and semantics; (2) the second cognitive process emphasizes Procedural Knowledge, training the model to plan and execute multi-step web interactions. To accommodate the increased complexity of the final stage, which involves multi-image inputs and extended reasoning, we configure training with a maximum sequence length of 8K and a batch size of 1 with gradient accumulation of 16 steps. All experiments are conducted on a cluster of 8 x NVIDIA A800 80GB GPUs.
5 Experiment
5.1 Experimental Setup
5.1.1 Models and Baselines
To comprehensively evaluate Web-CogReasoner, we compare its performance against several representative and state-of-the-art baselines.
Our proposed model, Web-CogReasoner, is built upon the Qwen2.5-VL-7Bbai2025qwen2 large multimodal model (LMM) as its foundational architecture. As detailed in Section 4, Web-CogReasoner extends Qwen2.5-VL by integrating a knowledge-induced cognitive reasoning framework, which leverages a structured knowledge base (described in Section 3) to guide chain-of-thought (CoT) reasoning for more robust and generalizable web navigation. Specifically, we performed supervised fine-tuning on Qwen2.5-VL using our curated datasets to imbue it with the proposed knowledge-driven capabilities.
[…]
These baselines collectively represent a diverse spectrum of current web agent methodologies, including foundational LLMs, interactive LLM-based agents, and powerful general-purpose LLM/LMMs leveraging prompting strategies. This allows for a thorough and multifaceted evaluation of Web-CogReasoner’s effectiveness and its unique advantages conferred by knowledge-induced cognitive reasoning.
5.1.2 Datasets
Training Dataset
In the training phase, we utilized a comprehensive cognitive dataset——Web-CogDataset,which is meticulously constructed as detailed in Section 3.3. This dataset is specifically designed to imbue Web-CogReasoner with knowledge-induced cognitive reasoning capabilities. It encompasses a rich variety of samples aligned with Factual, Conceptual, and Procedural knowledge dimensions, including diverse web page screenshots, corresponding accessibility trees, and detailed interaction trajectories with associated reasoning chains. Its comprehensive and structured nature is crucial for enabling the model to learn and perform sophisticated web navigation tasks.
Evaluation Datasets
To rigorously evaluate Web-CogReasoner’s performance and generalization capabilities, we employ a multi-faceted evaluation strategy utilizing various prominent datasets. We utilize the following datasets for evaluation:
  • Web-CogBench: As elaborated in Section 3.4, this custom-built benchmark is meticulously designed to assess distinct knowledge-driven cognitive abilities: Memorizing, Understanding, and Executing. Aallowing for granular evaluation of how our knowledge-induced reasoning framework contributes to these specific competencies.
  • VisualWebBench liu2024visualwebbench : A human-curated multimodal benchmark covering 1.5K examples across 139 real websites and 87 sub-domains, designed to evaluate MLLMs’s grounding and reasoning abilities in complex, web-based scenarios.
  • WebVoyager he2024openwebvoyager : This dataset comprises tasks on 15 websites that were seen during training and includes a total of 643 online web task queries. It primarily assesses the agent’s ability to perform tasks on familiar web environments.
  • Online Multimodal-Mind2Web deng2023mind2web : This includes two complementary subsets. The Cross-Task Test Set contains 112 queries from 33 websites seen during training, evaluating generalization to new tasks within known structures. The Cross-Website Test Set comprises 106 queries from 6 unseen websites across familiar domains (Entertainment, Shopping, and Travel), assessing the agent’s ability to transfer domain knowledge to novel layouts.
5.2 Evaluation Metrics
To provide a comprehensive and quantitative assessment of our model’s performance, we employ a set of specific metrics tailored to the nature of our evaluation datasets.
Web-CogBench Metrics
For our Web-CogBench, which evaluates distinct cognitive abilities, we use a variety of metrics depending on the task format:
  • Accuracy: For tasks with a definitive correct answer, such as multiple-choice questions (Next Page Prediction) or single-element identification (Source Element Prediction, Popup Close, Single Step Exploration), we calculate the standard accuracy. An answer is considered correct only if it exactly matches the ground truth.
  • ROUGE-L: For generative tasks that produce open-ended text, such as Element Attribute Recognition, we use ROUGE-L to measure the F1-score of the longest common subsequence between the predicted text and the ground truth. This metric effectively evaluates content overlap for short-form text generation.
  • LVM Judge: For tasks requiring qualitative assessment of generated content, such as Element Understanding and WebPage Understanding, a simple string match is insufficient. Therefore, we employ a powerful Large Vision Model (LVM) as an automated judge. The LVM is prompted to evaluate the quality, relevance, and accuracy of the model’s generated description against the ground truth, returning a numerical score on a 5-point scale. This approach provides a scalable and consistent method for evaluating complex, open-ended visual reasoning.
Online Web Tasks Metric
For online evaluation on benchmarks like WebVoyager and Multimodal-Mind2Web, the primary metric is the Task Success Rate (SR). A task is considered a success if and only if the agent successfully completes all required sub-steps and achieves the final goal specified in the instruction within the maximum allowed number of steps. The final success rate is the percentage of successfully completed tasks over the total number of tasks in the test set. This metric provides a clear and direct measure of the agent’s practical utility and effectiveness in real-world, dynamic environments.
5.3 Main Result
To validate the effectiveness of our proposed Web-CogReasoner and its Web-CogKnowledge Framework inspired by Bloom’s taxonomy, we conduct a comprehensive set of experiments across several challenging benchmarks. We evaluate our model’s performance in three key areas: fine-grained cognitive capabilities, visual understanding, and practical generalization on live online tasks. Our model is compared against leading proprietary agents, such as Claude Sonnet 4 and Gemini 2.5 Pro, as well as state-of-the-art open-source models.
Table 4: Performance evaluation on the Web-CogBench benchmark.
[…]
Overall
Claude Sonnet 4 54.3 64.7 100 96.8 76.8
Gemini 2.5 Pro 73.5 51.9 96.6 98.4 80.2
Qwen2.5-VL-7B 62.0 51.9 91.4 90.3 69.8
UI-TARs-7B-SFT 48.0 32.4 25.9 33.9 46.4
Web-CogReasoner (Ours) 79.0 61.4 98.3 95.2 84.4
Results on Web-CogBench
We first evaluate Web-CogReasoner on our proposed Web-CogBench, a benchmark specifically curated to assess the foundational knowledge and cognitive reasoning capabilities of web agents. It encompasses twelve diverse tasks ranging from element recognition to user intent inference, in alignment with our two-stage cognitive framework.
As presented in Table 4, Web-CogReasoner achieves state-of-the-art performance with an overall accuracy of 82.9%, surpassing both commercial and open-source baselines. Notably, it outperforms the powerful proprietary model Gemini 2.5 Pro (77.6%) and the strong open-source model Qwen2.5-VL-7B (66.9%), with substantial gains in high-level reasoning tasks such as Source Element Prediction (87.5% vs. 84.4%) and WebPage Understanding (79.0% vs. 73.5%). These results confirm the effectiveness of our design, which couples structured knowledge acquisition (factual, conceptual, procedural) with Knowledge-driven Chain-of-Thought (CoT) reasoning to form a solid cognitive foundation for agents.
Importantly, we find that strong performance on Web-CogBench is highly indicative of overall agent effectiveness across web tasks. For example, Gemini performs best on Web-CogBench among baseline models and also exhibits the best overall task completion rate in live-agent evaluations, in table 6 and table 7. However, we further observe that foundational model capabilities—particularly in visual perception—play a critical role in enabling such high-level cognition. Agents with weaker visual recognition skills exhibit reduced performance even when endowed with strong reasoning mechanisms, as reflected in additional evaluation results from VisualWebBench, in table 5.3, especially in perception-oriented tasks.
In summary, Web-CogBench serves not only as a direct measure of an agent’s reasoning capacity, but also implicitly requires strong perceptual grounding. Excelling on Web-CogBench implies that an agent possesses both the knowledge reasoning ability and the necessary perceptual competence—making it a highly reliable proxy for real-world web exploration effectiveness.
Table 5: Performance evaluation on the VisualWebBench benchmark. Tasks are grouped into Perception-Oriented and Reasoning-Oriented categories.
[…]
Ground
Claude Sonnet 4 73.3 72.6 96.2 80.7 81.1 96.1 96.3 91.2 85.9
Gemini 2.5 Pro 74.9 70.8 95.1 80.3 91.8 96.8 90.3 92.9 86.6
Qwen2.5-VL-7B 70.8 71.7 81.4 74.6 77.5 86.8 68.0 77.4 76.0
UI-TARs-7B-SFT 71.3 78.7 97.2 82.4 91.8 91.8 85.4 89.7 86.0
Web-CogReasoner (Ours) 67.2 72.6 97.0 79.0 96.4 96.1 88.4 93.6 86.3
Results on VisualWebBench
To evaluate the model’s visual understanding capabilities, we conduct experiments on the VisualWebBench benchmark liu2024visualwebbench . The results, summarized in Table 5.3, expose the limitations of visually-focused benchmarks and motivate the development of our proposed Web-CogBench.
Our Web-CogReasoner achieves the highest average score of 86.3%, marginally outperforming UI-TARs (86.0%). While this may suggest comparable performance at first glance, a closer examination reveals a stark contrast: UI-TARs performs significantly worse on Web-CogBench, with an overall score of only 48.2% (see Table 4).
This discrepancy underscores a critical insight — excelling in tasks such as OCR and element localization does not imply competence in cognitive reasoning. High scores on visual benchmarks like VisualWebBench can obscure deeper limitations in understanding and decision-making. In contrast, Web-CogReasoner demonstrates consistent strength across both visual and reasoning benchmarks, confirming that our framework effectively integrates visual perception with advanced cognitive reasoning. This dual capability is essential for building robust and reliable web agents.
Results on Online Web Tasks
Beyond offline benchmarks, we evaluated Web-CogReasoner’s practical utility and generalization ability on live, online web tasks using the WebVoyager he2024openwebvoyager and Online Multimodal-Mind2Web deng2023mind2web datasets. These experiments test the agent’s ability to apply its knowledge and reasoning to complete complex, multi-step and live tasks on unseen websites and scenarios, which is the ultimate goal for any autonomous web agent. Notably, we did not include UI-TARs-7B-SFT qin2025ui in the online evaluation due to the absence of publicly available online inference scripts. Therefore, we focus our comparison on the OpenWebVoyager he2024openwebvoyager series, which are specifically designed for web task execution.
Table 6: Task success rates on the WebVoyager test set. We compare our model, Web-CogReasoner, against several baselines across 14 websites. The "Overall" score is the average success rate.
Agent Allrecipes Amazon Apple ArXiv GitHub Booking ESPN Coursera
Claude Sonnet 4 26.7% 87.8% 48.8% 69.8% 68.3% 2.3% 45.5% 83.3%
Gemini 2.5 Pro 60.0% 63.4% 62.8% 67.4% 68.3% 9.1% 56.8% 73.8%
Qwen2.5-VL-7B 0.0% 0.0% 0.0% 4.7% 0.0% 0.0% 0.0% 2.3%
[…]
Overall Alpha
Claude Sonnet 4 23.8% 37.2% 4.8% 80.5% 48.8% 82.6% 47.7%
Gemini 2.5 Pro 52.3% 76.7% 4.8% 75.6% 58.1% 82.6% 54.9%
Qwen2.5-VL-7B 0.0% 11.6% 0.0% 2.4% 7.0% 2.2% 2.2%
[…]
Web-CogReasoner (Ours) 14.3% 55.8% 9.5% 39.0% 37.2% 39.1% 30.2%
[…]
Table 7: Performance comparison on the Mind2Web Online dataset under cross-task and cross-websites
Agent Cross-task (Unseen Task) Cross-web (Unseen Websites)
Claude Sonnet 4 44.9% 35.3% 40.0% 40.2% 45.5% 6.7% 14.0% 21.7%
Gemini 2.5 Pro 46.9% 35.3% 28.3% 37.5% 42.4% 10.0% 23.3% 25.5%
OpenWebVoyagerMax 22.4% 29.4% 15.2% 20.5% 3.0% 8.7% 23.3% 11.7%
Qwen2.5-VL-7B 2.2% 0.0% 0.0% 1.0% 3.0% 0.0% 0.0% 1.0%
OpenWebVoyagerIL 8.2% 5.9% 4.3% 6.3% 3.0% 5.8% 4.7% 6.6%
Web-CogReasoner (Ours) 16.3% 23.5% 15.2% 17.0% 12.1% 7.7% 9.3% 10.1%
The Mind2Web dataset (Table 7) serves as a crucial benchmark for evaluating generalization to unseen tasks (cross-task) and unseen websites (cross-web). Our model, Web-CogReasoner, records overall scores of 17.0% and 10.1% in these respective settings.
[…]
5.4 Ablation Study
Table 8: Ablation study of our Web-CogReasoner on the Web-CogBench. We evaluate the impact of each knowledge stage by progressively training the base model. "Base Model" is the original Qwen2.5-VL-7B. Each subsequent stage adds a new layer of knowledge training.
Model Configuration Memorizing Understanding Exploring Overall
Qwen2.5-VL-7B (Base Model) 67.6 61.0 77.9 69.8
+ Factual Knowledge (S1) 85.5 (+16.9) 64.2 60.1 72.1
+ Conceptual Knowledge (S2) 88.1 75.5 (+11.3) 65.8 78.3
+ Procedural Knowledge (S3) 90.8 74.1 85.0 (+7.1) 84.4
To empirically validate the effectiveness of our hierarchical, Bloom’s Taxonomy-inspired training methodology, we conduct a thorough ablation study. Our goal is to isolate and quantify the contribution of each distinct knowledge layer: Factual, Conceptual, and Procedural. We progressively build our Web-CogReasoner agent, starting from the base Qwen2.5-VL-7B model and sequentially adding each knowledge stage. The performance at each step is evaluated on our Web-CogBench, as its cognitive dimensions (Memorizing, Understanding, Exploring) directly align with our knowledge framework. The results are presented in Table 8.
Impact of Factual Knowledge
The first stage of our curriculum introduces Factual Knowledge, teaching the agent the "what" of web elements. As shown in Table 8, this initial stage provides a substantial performance boost, raising the overall score from 66.9% to 72.1%. Critically, the improvement is most pronounced in the Memorizing dimension, which surges from 68.6% to 85.5%. This demonstrates that by learning to perceive and identify concrete, factual information (e.g., element roles and names), the agent builds a solid perceptual foundation, which is the first and essential step towards more complex reasoning.
Impact of Conceptual Knowledge
Building upon this factual base, the second stage integrates Conceptual Knowledge, which focuses on the "how things relate." This layer enables the agent to understand the relationships between elements and the underlying semantic structure of a webpage. The results show another significant leap in performance, with the overall score climbing to 78.3%. The most notable gain occurs in the Understanding dimension, which improves by 11.3%. This confirms our hypothesis that once an agent can perceive facts, it must then learn to comprehend their context and purpose to effectively interpret a webpage’s functionality.
Impact of Procedural Knowledge
The final stage incorporates Procedural Knowledge, teaching the agent "how to act" to achieve a goal. This layer completes the Web-CogReasoner’s cognitive architecture, enabling it to formulate and execute plans. This stage delivers the final and most critical improvements for task completion, pushing the overall score to our final state-of-the-art result of 85.0%. The most dramatic increase is observed in the Exploring dimension, which skyrockets by 7.1%. This finding strongly indicates that goal-oriented, procedural training is paramount for converting passive understanding into active, successful task execution.
In summary, this staged analysis provides compelling empirical evidence for our core thesis. Each knowledge layer, from Factual to Conceptual to Procedural, is an indispensable component that progressively builds the agent’s cognitive capabilities, validating our structured, curriculum-based approach to training highly capable web agents.
5.5 Qualitative Analysis
Beyond the quantitative improvements shown in our ablation study, a qualitative analysis of the agent’s behavior at each stage offers deeper insights into how our curriculum shapes its cognitive abilities. We examine the agents’ performance on a representative task: "Find and add a laptop under \$1000 to the shopping cart on an e-commerce website."
Base Model (Qwen2.5-VL-7B)
Without any specialized training, the base model struggles to formulate a coherent plan. Its reasoning is often generic and untethered from the specific UI. It might correctly identify a "search bar" but fails to execute a meaningful action, or hallucinates actions that are not possible. For instance, its thought process might be: "I should search for a laptop," but its action is an ungrounded ‘click "Categories"‘ because it lacks the procedural knowledge to connect intent to a multi-step sequence of actions.
Stage 1 Agent (+ Factual Knowledge)
After training on Factual Knowledge, the agent’s perceptual abilities are significantly enhanced. It can now accurately identify and label key elements with their correct attributes. Its thought process becomes grounded in the facts of the page: "I see a search bar [ID: 25] with the name ’Search products’. I see a button [ID: 28] with the name ’Search’." However, it still struggles with planning. It understands "what" is on the page but not "why" or "how" to use it. It might correctly type "laptop" into the search bar but then get stuck, not understanding that the next logical step is to click the search button to submit the query.
Stage 2 Agent (+ Conceptual Knowledge)
With the addition of Conceptual Knowledge, the agent begins to understand the relationships between elements and their purpose. Its reasoning graduates from simple identification to semantic interpretation. The thought process now reflects this understanding: "The search bar [ID: 25] is for inputting queries. The search button [ID: 28] is functionally linked to it and will trigger the search. This group of elements forms a ’search component’." This allows it to reliably complete the search and navigate to the results page. However, on the results page, it may still struggle with complex procedural logic, such as applying a price filter.
Full Model (Web-CogReasoner, + Procedural Knowledge)
The final agent, equipped with Procedural Knowledge, demonstrates goal-oriented planning and execution. It seamlessly translates the high-level task into a concrete action sequence. Its thought process is now a strategic plan: "Goal: Add laptop under \$1000 to cart. Step 1: Type ’laptop’ into search bar [ID: 25]. Step 2: Click search button [ID: 28]. Step 3: On the results page, locate the ’Price Range’ filter. Step 4: Input ’1000’ into the ’max price’ field [ID: 57]. Step 5: Identify a suitable product from the filtered list and click its ’Add to Cart’ button [ID: 83]." This demonstrates a complete cognitive loop from perception and understanding to successful action, validating the necessity of the final procedural training stage.
Appendix A Appendix
A.1 Bloom’s Taxonomy and Web Cognitive
A.1.1 Bloom’s Tax
  • 1.
  • 2.
  • 3.
    Procedural Knowledge: The knowledge of how to perform a task or inquiry. This involves an understanding of specific skills, algorithms, techniques, and methods, representing a shift from "knowing-what" to "knowing-how."
A.1.2 Web data introduction
A.1.3 Data Sourcing
@@ A.2 Task definition
A.2.1 Factual Web Knowledge
Element Attribute Rec
@@ A.2.1 Factual Web Knowledge
  • the semantic role of the highlighted element (e.g., "button", "link", "checkbox"),
  • the semantic name, which refers to the element’s accessible textual description (e.g., "Submit", "Search", "Next").
@@
Page Change Pre
We define the Page Change Prediction task to evaluate a model’s capability to infer the visual consequences of interacting with a specific web element, relying solely on visual input. In this task, the model is presented with a full-page screenshot in which a target interactive element is highlighted by a red bounding box. The model is required to in an open-ended format to predict the visual changes that are likely to occur on the page after the element is clicked.The ground truth for this task is obtained from the generated responses of Qwen-VL 72B, which were produced based on visual metadata, as detailed in SectionA.1.3. This task is designed to simulate the human cognitive ability to anticipate the dynamic behavior of web interfaces through perception alone—without access to the underlying source code or prior knowledge of the page logic.
Next Page Pre

### v2

@@ A.1.4 Web data introduction
    Breadth via Open-Source Augmentation: To address the concern of limited domain diversity (e.g., lack of News, Education, or Forums), we incorporated and enhanced large-scale open-source datasets, including MultiUI (Liu et al., 2024a), Mind2Web (Deng et al., 2023), and OpenWebVoyager (He et al., 2024b). Notably, MultiUI is derived from FineWeb (Common Crawl), providing massive coverage of general-purpose webpages. This combination ensures our model generalizes to the ”wild” web and is not overfitted to specific interaction styles like financial trading or shopping.
Annotation Reliability
We validated our automated annotations via double-blind human verification and cross-model consistency checks (e.g., using GPT-4o). As shown in Table 11, the error rate is minimal.
Table 11: Reliability Check of Web-CogDataset Annotations.
Annotation Task Human Verification (Acc) Cross-Model Consistency
Element Attribute 99.2% 98.5%
Page Change Pred 97.5% 96.8%
Sub-element Pred 96.8% 95.4%
Average 97.8% 96.9%
A.1.5 Data Sourcing
@@ A.2 Task definition
A.2.1 Web-CogDataset
Table 13: Statistics of Web-CogDataset.
Knowledge Task Statistics
Subtotal Total
Factual Web Knowledge Element Attribute Recognition 37K 81K
Sub-element Prediction 1K
Page Change Prediction 24K
Next Page Prediction 18K
Source Element Prediction 1K
Conceptual Web Knowledge Element Understanding 40K 62K
WebPage Understanding 7K
Caption & QA 15K
Procedural Web Knowledge User’s Intention Prediction 2K 27K
Popup Close 1K
Single-Step Web Task 17K
Noisy Multi-Step Web Task 7K
A.2.2 Web-CogBench
  • Element Attribute Recognition: Given a screenshot with a highlighted interactive element, the model predicts its semantic role (e.g., button, link) and accessible name (e.g., ”Submit”, ”Search”), relying solely on visual cues.
  • Next Page Prediction: The model predicts the subsequent page that results from interacting with a specific element on the current page. To enhance generalization, we designs two types of tasks: multiple-choice questions and open-ended responses.
  • Source Element Prediction: Given two screenshots, the current page and the resulting target page, then the model identifies which of the visually marked elements on the current page leads to the target, simulating visual cause-and-effect reasoning.
  • User’s Intention Prediction: The model infers high-level user intent from a sequence of webpage screenshots representing an interaction trajectory, requiring visual understanding and temporal reasoning. The task is built on the MultiModal-Mind2Web (Deng et al., 2023) dataset, mapping screenshot sequences to natural language instructions.
  • Popup Close: The model identifies and dismisses popups (e.g., notification modals, login forms) on synthesized webpage screenshots, using a dataset of 51 popup components from JS Design [^3] https://js.design/. overlaid on OpenWebVoyager (He et al., 2024b) webpages, with combinatorial augmentation of closing strategies for diverse training.
  • Single Step Exploration: This task is derived from a multi-step trajectory exploration task and has been decomposed into single-step exploration subtasks. For each step, the model receives as input the corresponding accessibility tree (AxTree) and a screenshot of the current webpage. Based on these observations, the model performs reasoning to generate the appropriate action along with the target object, effectively simulating a realistic single-step web navigation scenario. By breaking down complex multi-step interactions into atomic actions, this setup allows for fine-grained evaluation of the agent’s decision-making capabilities and supports systematic analysis of its performance in web exploration tasks.
Table 13 summarizes the overall statistics of Web-CogDataset. The dataset covers three layers of knowledge—Factual, Conceptual, and Procedural— each corresponding to distinct families of web reasoning tasks. Factual Web Knowledge focuses on recognizing attributes, predicting element relationships, and modeling page transitions, totaling 81K instances. Conceptual Web Knowledge emphasizes semantic understanding and cross-element comprehension, with 62K instances. Procedural Web Knowledge involves action-oriented reasoning tasks, such as predicting user intentions and executing goal-directed interactions, comprising 27K instances. Together, these task distributions reflect the hierarchical design of Web-CogDataset and ensure balanced coverage from low-level perception to high-level reasoning.
A.2.3 Factual Web Knowledge
Element Attribute Recognition
@@ A.2.3 Factual Web Knowledge
  • the semantic role of the highlighted element (e.g., ”button”, ”link”, ”checkbox”),
  • the semantic name, which refers to the element’s accessible textual description (e.g., ”Submit”, ”Search”, ”Next”).
@@
Page Change Prediction
We define the Page Change Prediction task to evaluate a model’s capability to infer the visual consequences of interacting with a specific web element, relying solely on visual input. In this task, the model is presented with a full-page screenshot in which a target interactive element is highlighted by a red bounding box. The model is required to in an open-ended format to predict the visual changes that are likely to occur on the page after the element is clicked.The ground truth for this task is obtained from the generated responses of Qwen-VL-72B, which were produced based on visual metadata, as detailed in SectionA.1.5. This task is designed to simulate the human cognitive ability to anticipate the dynamic behavior of web interfaces through perception alone—without access to the underlying source code or prior knowledge of the page logic.
Next Page Prediction